if i have the choice of the following methods of payment how much would i have to earn in sales to get the same pay regardless of the option chosen and when would i be better off in option A?
Payment Method A: A retainer of $250 per week plus 2.5% of sales.
Payment Method B: No retainer but 3.5% of sales.

Answer by Theo(13342)

method A.

250 per week retainer plus 2.5% of sales.

method B.

no retainer plus 3.5% of sales.


equation for method A.

r = 250 + .025 * x

equation for method B.

r = .035 * x

when the revenues are equal, 250 + .025 * x = .035 * x

subtract .025 * x from both sides of the equation to get 250 = .01 * x

solve for x to get x = 250 / .01 = 25000.

the two methods break even when the store makes sales equal to 25000.

when that happens, r = 250 + .025 * 25000 = 875 and r = .035 * 25000 = 875.

you would be better off with method A when the average amount of sales per week is less than 25000.